The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 64 June 23, 2002

Contents

Letter From The Editor  1
Abstracts of 23 refereed papers  2
Abstracts of 2 conference contributions 17
Abstract of 1 thesis  19

Letter From The Editor

Dear Colleagues,

Due to increased time constraints, I can no longer act as editor-in-chief for the Magellanic Clouds Newsletter. This has led to a lengthy delay in the publication of the current issue, which I regret. Fortunately, You-Hua and I have found a person excellently suited to take over the duties of the primary editor: Bryan Dunne (carolan@astro.uiuc.edu) at the University of Illinois in Urbana-Champaign. Bryan has taken over as editor starting with the current issue, while You-Hua and I will remain coeditors. With Bryan on board we are confident that MCNews will return to its monthly publication schedule. So please keep supplying Bryan with a steady stream of contributions. The e-mail address of MCNews as well as its URLs will remain the same. We hope to provide on-line abstract submission soon.

Many thanks to Bryan, and best of luck!

Best wishes, Eva Grebel & You-Hua Chu

Abstracts of Refereed Papers

An ATCA Radio-Continuum Study of the Small Magellanic Cloud Part 1: Source Catalogues at 1.42, 2.37, 4.80 and 8.64 GHz

M. D. Filipović,^{1,2} T. Bohlsen,^{1} W. Reid,^{1} L. Staveley-Smith,^{2} P. A. Jones^{1}, K. Nohejl^{1} and G. Goldstein^{1}

1University of Western Sydney, Locked Bag 1797, Penrith South, DC, NSW 1797, Australia 2Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia

We have analysed radio-continuum mosaics of the Small Magellanic Cloud (SMC) obtained using both the Australia Telescope Compact Array (ATCA) and the Parkes telescope and present a new catalogue of sources in the SMC at 1.42, 2.37, 4.80 and 8.64 GHz (λ=20, 13, 6 and 3 cm). We find a total of 717 radio-continuum sources at these frequencies in the field of the SMC which is three times more than previous Parkes surveys. From these 717 radio-continuum sources, some 534 were detected at 1.42 GHz, 697 at 2.37 GHz, 75 at 4.80 Ghz and 54 at 8.64 GHz. The integrated flux density was measured for each of these sources. We have assessed the accuracy of the position and flux measurements of our catalogue and found no significant discrepancy with previous catalogues. Our new data shows a significant improvement in sensitivity (over 10 times) and positional accuracy (< 1′′) over previous catalogues.

Accepted by: MNRAS

For preprints, contact M. D. Filipović your e-mail address m.filipovic@uws.edu.au

ROSAT PSPC view of the hot interstellar medium of the Magellanic Clouds

M. Sasaki, F. Haberl, and W. Pietsch

Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, Postfach 1312, 85741 Garching, Germany

Diffuse X-ray emission from the Magellanic Clouds (MCs) is studied by using all the archival data of pointed ROSAT Position Sensitive Proportional Counter (PSPC) observations. For this purpose, contributions from the point and point-like sources in the ROSAT High Resolution Imager (HRI) and PSPC source catalogues are eliminated and periods of high solar activity are excluded. The spectral analysis yielded characteristic temperatures of 10^{6} − 10^{7} K for the hot thin plasma of the ISM which extends over the whole Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC). The total unabsorbed luminosity in the 0.1 – 2.4 keV band within the observed area amounts to 3.2 × 10^{38} erg s^{−1} in the LMC and 1.1 × 10^{37} erg s^{−1} in the SMC, each with an uncertainty of ∼ −40%, +100%. The X-ray luminosity of the LMC is comparable to that of other nearby galaxies with pronounced star formation. In the LMC, hot regions were found especially around the supergiant shell (SGS) LMC 4 and in the field covering SGS LMC 2 and LMC 3. Highest temperatures for the SMC were derived in the southwestern part of the galaxy. The diffuse X-ray emission is most likely a superposition of the emission from the hot gas in the interior of shells and supershells as well as from the halo of these galaxies.

Accepted by: Astronomy & Astrophysics For preprints, contact manami@mpe.mpg.de

Also available from the URL http://arXiv.org/abs/astro-ph/0206335

Further Insights into the Structure of 30 Doradus from the Hubble Space Telescope Instruments

Nolan R. Walborn^{1}, Jesús Maíz-Apellániz^{1}, and Rodolfo H. Barbá^{2}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, Paseo del Bosque, 1900 La Plata, Argentina

New observations with the HST Wide Field Planetary Camera 2, combined with prior archival data, provide nearly complete coverage of the inner 30 Doradus Nebula in both nebular lines and continuum at 0.′′1 resolution. The developing wind-blown cavity surrounding the massive first-generation, central cluster and its interface with the remanent molecular clouds, the site of second-generation, triggered star formation, are imaged in their entirety. Dithered Planetary Camera observations with 0.′′03 resolution of Knots 1–3 reveal further structural details of the large dust pillars oriented toward the central cluster and the newborn, massive multiple systems of the second generation in those fields. The new data also provide the first WFPC2 coverage of the fields of two interesting, luminous infrared sources observed in our previous Near Infrared Camera and Multi-Object Spectrometer program; comparisons of the high-resolution optical and IR images are illuminating. In addition, we have obtained Space Telescope Imaging Spectrograph optical, long-slit observations of seven O-type multiple systems in the region, including those in Knots 1–3 and four representatives of the older generation. The spatially resolved spectrograms and classifications of the close pairs, with separations ranging from 0.′′09 to 1.′′18, are presented, as well as WFPC2 photometry of the individual components and the compact clusters of fainter stars associated with them. These new observations, and planned further analysis of them, offer significant new information about the intricate structure and evolution of the two-stage starburst in 30 Doradus.

Accepted by The Astronomical Journal

Preprints from walborn@stsci.edu

Modelling the Hubble Space Telescope Ultraviolet and Optical Spectrum of Spot 1 on the Circumstellar Ring of SN 1987A

C. S. J. Pun^{1,2,3} and the Supernova INtensive Study (SINS) team

1 Laboratory for Astronomy and Space Physics, Code 681, NASA-GSFC, Greenbelt, MD 20771, USA 2 National Optical Astronomical Observatories, P.O. Box 26732, Tucson, AZ 85726, USA 3 Present address: Dept. of Physics, University of Hong Kong, Pokfulam Road, Hong Kong

We report and interpret HST/Space Telescope Imaging Spectrograph (STIS) long-slit observations of the optical and ultraviolet (UV) (1150 − 10270 Å) emission-line spectra of the rapidly brightening Spot 1 on the equatorial ring of SN 1987A between 1997 September and 1999 October (days 3869 – 4606 after outburst). The emission is caused by radiative shocks created where the supernova blast wave strikes dense gas protruding inward from the equatorial ring. We measure and tabulate line identifications, fluxes and, in some cases, line widths and shifts. We compute flux correction factors to account for substantial interstellar line absorption of several emission lines. Nebular analysis shows that optical emission lines come from a region of cool (T_{e} ≈ 10^{4} K) and dense (n_{e} ≈ 10^{6}cm^{−3}) gas in the compressed photoionized layer behind the radiative shock. The observed line widths indicate that only shocks with shock velocities V_{s} < 250 km s^{−1} have become radiative, while line ratios indicate that much of the emission must have come from yet slower (V_{s} < ∼ 135 km s^{−1}) shocks. Such slow shocks can be present only if the protrusion has atomic density n > ∼ 3 × 10^{4}cm^{−3}, somewhat higher than that of the circumstellar ring. We are able to fit the UV fluxes with an idealized radiative shock model consisting of two shocks (V_{s} = 135 and 250 km s^{−1}). The observed UV flux increase with time can be explained by the increase in shock surface areas as the blast wave overtakes more of the protrusion. The observed flux ratios of optical to highly-ionized UV lines are greater by a factor of ∼ 2 − 3 than predictions from the radiative shock models and we discuss the possible causes. We also present models for the observed Hα line widths and profiles, which suggests that a chaotic flow exists in the photoionized regions of these shocks. We discuss what can be learned with future observations of all the spots present on the equatorial ring.

Accepted by: The Astrophysical Journal (Part I) For preprints, contact jcspun@hkucc.hku.hk

Stellar and Wind Properties of LMC WC4 stars A metallicity dependence for Wolf-Rayet mass-loss rates

Paul A. Crowther (1), Luc Dessart (2), D. John Hillier (3), Jay B. Abbott (1), A. W. Fullerton (4)

(1) Dept. of Physics & Astronomy, UCL, Gower St, London WC1E 6BT, U.K.;

(2) Sterrenkundig Institut, Universiteit Utrecht, P.O. Box 80000, NL-3508 TA Utrecht, The Netherlands;

(3) Department of Physics & Astronomy, University of Pittsburgh, 3941 O’Hara Street, PA 15260, USA;

(4) Department of Physics & Astronomy, Johns Hopkins University, 3400 North Charles St., Baltimore, MD 21218, USA

We use ultraviolet space-based (FUSE, HST) and optical/IR ground-based (2.3m MSSSO, NTT) spectroscopy to determine the physical parameters of six WC4-type Wolf-Rayet stars in the Large Magellanic Cloud. Stellar parameters are revised significantly relative to Gräfener et al. (1998) based on improved observations and more sophisticated model atmosphere codes, which account for line blanketing and clumping. We find that stellar luminosities are revised upwards by up to 0.4 dex, with surface abundances spanning a lower range of 0.1≤ C/He ≤ 0.35 (20–45% carbon by mass) and O/He≤0.06 (≤10% oxygen by mass).

Relative to Galactic WC5–8 stars at known distance analysed in a similar manner, LMC WC4 stars possess systematically higher stellar luminosities, ∼0.2 dex lower wind densities, yet a similar range of surface chemistries. We illustrate how the classification C iii λ5696 line is extremely sensitive to wind density, such that this is the principal difference between the subtype distribution of LMC and Galactic early-type WC stars. Temperature differences do play a role, but carbon abundance do not affect WC spectral types. We illustrate the effect of varying temperature and mass-loss rate on the WC spectral type for HD 32257 (WC4, LMC) and HD 156385 (WC7, Galaxy) which possess similar abundances and luminosities.

We determine current stellar masses of LMC WC4 stars, taking account of the (minor) effect of wind darkening, once clumped winds are considered. Depending on the evolutionary model followed, pre-supernova masses 11–19 M_{⊙} are anticipated, with 7–14 M_{⊙} for Galactic WC stars with known distances. These values are consistent with pre-cursors of bright Type-Ic supernovae such as SN 1998bw (alias GRB 980425) for which a minimum total mass of C and O of 14M_{⊙} has been independently derived.

Accepted by: Astronomy & Astrophysics For preprints, contact pac@star.ucl.ac.uk

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0206233

Revised Stellar Temperatures for Magellanic Cloud O Supergiants from FUSE and VLT-UVES Spectroscopy

Paul A. Crowther (1), D.J. Hillier (2), C.J. Evans (1), A. W. Fullerton (3), O. De Marco (4), A.J. Willis (1)

(1) Dept. of Physics & Astronomy, UCL, Gower St, London WC1E 6BT, U.K.;

(2) Department of Physics & Astronomy, University of Pittsburgh, 3941 O’Hara Street, PA 15260, USA;

(3) Department of Physics & Astronomy, Johns Hopkins University, 3400 North Charles St., Baltimore, MD 21218, USA;

(4) Department of Astrophysics, American Museum of Natural History, Central Park West at 79th St, New York, NY 10024, USA

We have undertaken quantitative analysis of four LMC and SMC O4–9.7 extreme supergiants using far-ultraviolet FUSE, ultraviolet IUE/HST and optical VLT UVES spectroscopy. Extended, non-LTE model atmospheres that allow for the consistent treatment of line blanketing (Hillier & Miller 1998) are used to analyse wind and photospheric spectral features simultaneously. Using Hα to constrain Ṁ, He i-ii photospheric lines reveal stellar temperatures which are systematically (5–7.5kK) and substantially (15–20%) lower than previously derived from unblanketed, plane-parallel, non-LTE photospheric studies. We have confidence in these revisions, since derived temperatures generally yield consistent fits across the entire λλ912–7000Å observed spectral range. In particular, we are able to resolve the UV-optical temperature discrepancy identified for AzV232 (O7 Iaf^{+}) in the SMC by Fullerton et al. (2000).

The temperature and abundance sensitivity of far-UV, UV and optical lines is discussed. ‘Of’ classification criteria are directly linked to (strong) nitrogen enrichment (via N iii λ4097) and (weak) carbon depletion (via C iii λλ4647-51), providing evidence for mixing of unprocessed and CNO processed material at their stellar surfaces. Oxygen abundances are more difficult to constrain, except via O ii lines in the O9.7 supergiant for which it is also found to be somewhat depleted. Unfortunately, He/H is very difficult to determine in individual O supergiants, due to uncertainties in microturbulence and the atmospheric scale height. The effect of wind clumping is also investigated, for which P v λλ1118– 28 potentially provides a useful diagnostic in O-star winds, unless phosphorus can be independently demonstrated to be underabundant relative to other heavy elements. Revised stellar properties affect existing calibrations of (i) Lyman continuum photons – a factor of two lower for the O4 supergiant; and (ii) kinetic energy released into the ISM by O supergiants. Our results also have importance for the calibration of the wind momentum-luminosity relationship for OB stars, particularly since the stars studied here are amongst the visually brightest OB stars in external galaxies.

Accepted by: Astrophysical Journal For preprints, contact pac@star.ucl.ac.uk

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0206257

Far Ultraviolet Spectroscopic Explorer Atlas of OB Stars in the Magellanic Clouds

Nolan R. Walborn^{1}, Alex W. Fullerton^{2,3}, Paul A. Crowther^{4}, Luciana Bianchi^{3,5}, John B. Hutchings^{6}, Anne Pellerin^{7}, George Sonneborn^{8}, and Allan J. Willis^{4}

1Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2Department of Physics & Astronomy, University of Victoria, P.O. Box 3055, Victoria, BC, V8W 3P6, Canada 3Department of Physics & Astronomy, The Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 4Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK 5On leave from the Astronomical Observatory of Torino 6Herzberg Institute of Astrophysics, National Research Council of Canada, 5071 West Saanich Road, Victoria, BC, V8X 4M6, Canada 7Département de Physique, de Génie Physique, et d’Optique, Université Laval, Québec, QC, G1K 7P4, Canada 8Laboratory for Astronomy & Solar Physics, NASA Goddard Space Flight Center, Code 681, Greenbelt, MD 20771, USA

An atlas of the 900–1200 Å region in the spectra of 47 OB stars in the Large and Small Magellanic Clouds, observed at high resolution by the Far Ultraviolet Spectroscopic Explorer, is presented and discussed. The systematic trends in the numerous stellar-wind features in this region, some from species (and ionizations) not represented at longer wavelengths, are charted as a function of the optical spectral types. The FUSE sample is by far the most powerful to date for that purpose. A special effort has been made to verify the spectral types of all stars included in the atlas, in a number of cases with new optical observations that are also illustrated, to avoid uncertainties from that source in the stellar-wind trends. A new O2 star has been found in the process. Most of these stars have been previously observed at longer ultraviolet wavelengths by the Hubble Space Telescope, and in the optical from the ground with high-resolution, digital instruments; thus very comprehensive physical modeling of these OB atmospheres and winds now becomes possible. This atlas will serve as a guide to the FUSE Magellanic Cloud OB database for that purpose. The Magellanic Cloud sample provides a very important complement to the FUSE database of Galactic OB counterparts (Pellerin et al. 2002), both because the lower extinction and interstellar H_{2} absorption toward the Cloud stars allow a much clearer view of the stellar spectra below 1100 Å, and because of the metallicity differences among the three galaxies. In particular, most wind features in the SMC spectra are seen to be significantly weaker than those in the LMC, at the same spectral types.

Accepted by Astrophysical Journal Supplements

Preprints from walborn@stsci.edu

Massive luminous early type stars in the LMC

I. The reddening of individual stars and the LMC reddening law

J. Gochermann^{1,2,3} and Th. Schmidt-Kaler^{2,4}

1 Institut für Technische Betriebswirtschaft, Fachhochschule Münster, Stegerwaldstraße 39, D-48565 Steinfurt, Germany 2 Astronomisches Institut der Ruhr-Universität Bochum, D-44780 Bochum, Germany 3 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, D-85748 Garching, Germany 4 Georg-Büchner-Straße 37, D-97276 Margetshöchheim, Germany

In order to construct a comprehensive HRD of early type stars in the Large Magellanic Cloud (LMC) in the first step the reddenings of individual stars and the LMC reddening law have been investigated. 1942 LMC member stars with good UBV photometries from the Bochum photometry data base have been first corrected individually for galactic foreground reddening. From stars with good spectral classification the slope of the LMC internal reddening line (\frac{E_{U−B}}{E_{B−V}} )was calculated for each spectral subclass between O3 and A4. A remarkable difference to the galactic reddening law was found. The slope of the reddening line first decreases for stars from O3 to B0, and then increases rapidly between B0 and B3 from ∼ 0.7 to ∼ 1.1. For later type stars it remains higher than for early type stars. This effect has important consequences for all extinction corrections. We checked this using different methods. Because no evidence for systematically wrong classifications was found, the differences in the reddening slopes must be caused by the ISM of the LMC itself. Four possible causes are considered.

Accepted by: Astronomy & Astrophysics

For preprints, contact gochermann@fh-muenster.de

Also available from the URL

http://www.astro.ruhr-uni-bochum.de/astro/publications/pub2002.html

Discovery of a Possible Anomalous X-ray Pulsar in the Small Magellanic Cloud

R.C. Lamb, D.W. Fox, T.A. Prince^{1}, and D.J. Macomb^{2}

1 Space Radiation Laboratory, California Institute of Technology, Pasadena, CA 91125, USA 2 Physics Department, Boise State University, Boise, ID 83725-1570, USA

We report the serendipitous detection of a previously unreported pulsar from the direction of the Small Magellanic Cloud using data from the CHANDRA X-Ray Observatory. Because of its luminosity (∼ 1.5 × 10^{35} ergs/s), its near lack of variability for more than 20 years, and its very soft spectrum we propose that it is an anomalous X-ray pulsar (AXP). Data from the ROSAT PSPC in conjunction with the CHANDRA data give a period, P, of 5.44 s and a spin down time, P/Ṗ, of 11 ky. If this is a correct identification it will be the first extragalactic AXP and the fastest yet discovered.

Accepted by: Astrophysical Journal Letters

For preprints, contact lamb@srl.caltech.edu

Supernova 1987A did not test the neutrino mass hierarchy

1V. Barger, ^{2}D. Marfatia and ^{1}B. P. Wood

1 Department of Physics, University of Wisconsin, Madison, WI 53706, USA 2 Department of Physics, Boston University, Boston, MA 02215, USA

We dispel the misconception that data from SN 1987A favor the normal neutrino mass hierarchy over the inverted hierarchy for sin^{2} θ_{13} > ∼ 10^{−4}. We find comparable fits for the two hierarchies. No bound can be placed on the mixing angle θ_{13} even at the 1σ level.

Accepted by: Physics Letters B.

For preprints, contact marfatia@physics.bu.edu
Also available from the URL http://arXiv.org/abs/hep-ph/0202158

The Massive Wolf-Rayet Binary SMC WR7

V.S. Niemela ^{1}, P. Massey ^{2}, G. Testor ^{3} and S. Giménez Benítez ^{1}

1Facultad de Cs. Astronómicas y Geofísicas, Universidad Nacional de La Plata, Paseo del Bosque s/n, 1900 La Plata, Argentina; 2 Lowell Observatory, 1400 West Mars Hill Road, Flagstaff, AZ 86001, USA; 3 Observatoire de Paris, section de Meudon, F-92195 Meudon Cedex, DAEC, France;

We present a study of optical spectra of the Wolf–Rayet star AzV 336a (= SMC WR7) in the Small Magellanic Cloud. Our study is based on data obtained at several Observatories between 1988 and 2001. We find SMC WR7 to be a double lined WN+O6 spectroscopic binary with an orbital period of 19.56 days. The radial velocities of the He absorption lines of the O6 component and the strong Heii emission at λ4686Å of the WN component describe antiphased orbital motions. However, they show a small phase shift of ∼ 1 day. We discuss possible explanations for this phase shift. The amplitude of the radial velocity variations of He ii emission is twice that of the absorption lines. The binary components have fairly high minimum masses, ∼ 18 M_{⊙} and 34 M_{⊙} for the WN and O6 components, respectively.

Accepted by: MNRAS

For preprints, contact virpi@fcaglp.unlp.edu.ar
Also available from the URL http://lilen.fcaglp.unlp.edu.ar/papers2002.html

Large Magellanic Cloud Bump Cepheids: Probing the Stellar Mass-Luminosity Relation

Stefan C. Keller ^{1} & Peter R. Wood ^{2}

1 L-413, IGPP, LLNL, PO Box 505 Livermore, CA 94550, USA 2 RSAA, Australian Nat. Univ., Canberra ACT 2600, AUSTRALIA

We present the results of non-linear pulsation modelling of 20 bump Cepheids in the LMC. By obtaining a optimal fit to the observed V,R MACHO lightcurves we have placed tight constraints on stellar parameters of M, L, T_{eff} and well as quantities of distance modulus and reddening. We describe the mass-luminosity relation for core-He burning for intermediate mass stars. The mass-luminosity relation depends critically on the level of mixing within the stellar interior over the course of the main-sequence lifetime. Our sample is significantly more luminous than predicted by classical stellar evolutionary models that do not incorporate extension to the convective core. Under the paradigm of convective core overshoot our data implies Λ_{c} of 0.65±0.03l/H_{p}. We derive a LMC distance modulus of 18.55±0.02.

Accepted by: Astrophysical Journal

For preprints, contact skeller@igpp.ucllnl.org

Also available from the URL astro-ph 0205555

Early-type variables in the Magellanic Clouds. I. β Cephei stars in the LMC bar

A. Pigulski and Z. Kołaczkowski

Wrocław University Observatory, Kopernika 11, 51-622 Wrocław, Poland

A thorough analysis of the OGLE-II time-series photometry of the Large Magellanic Cloud bar supplemented by similar data from the MACHO database led us to the discovery of three β Cepheitype stars. These are the first known extragalactic β Cephei-type stars. Two of the three stars are multiperiodic. Two stars have inferred masses of about 10 M_{⊙} while the third is about 2 mag brighter and at least twice as massive. All three variables are located in or very close to the massive and young LMC associations (LH 41, 59 and 81). It is therefore very probable that the variables have higher than average metallicities. This would reconcile our finding with theoretical predictions of the shape and location of the β Cephei instability strip in the H-R diagram. The low number of β Cephei stars found in the LMC is another observational confirmation of strong dependence of the mechanism driving pulsations in these variables on metallicity. Follow-up spectroscopic determination of the metallicities in the discovered variables will provide a good test for the theory of pulsational stability in massive main-sequence stars.

Accepted by: Astronomy & Astrophysics appear

For preprints, contact pigulski@astro.uni.wroc.pl
Also available from the URL http://www.astro.uni.wroc.pl/nauka/papers/h3487.ps.gz

Convective Core Mixing: a Metallicity Dependence?

D. Cordier^{1,2}, Y. Lebreton^{1}, M.-J. Goupil^{1}, T. Lejeune^{3}, J.-P. Beaulieu^{4} and F. Arenou^{1}

1 Observatoire de Paris-Meudon, DASGAL, F-92195 Meudon Principal Cedex, France. 2 École Nationale Supérieure de Chimie de Rennes, Campus de Beaulieu, F-35700 Rennes, France 3 Observatorio Astronomico, Universidade de Coimbra, Santa Clara 3040 Coimbra, Portugal. 4 I.A.P., 98bis boulevard Arago, F-75014 Paris, France.

The main scope of this paper is to investigate the possible existence of a metallicity dependence of the overshooting from main sequence stars turbulent cores. We focus on objects with masses in the range ∼ 2.5 M_{⊙}− ∼ 25 M_{⊙}. Basically, evolutionary time scale ratios are compared with star numbers ratios on the main sequence. Star populations are synthesized using grids of evolutionary tracks computed with various overshooting amounts. Observational material is provided by the large and homogeneous photometric database of OGLE 2 project for the Magellanic clouds. Attention is paid to the study of uncertainties: distance modulus, intergalactic and interstellar reddening, IMF slope and average binarity rate. Rotation and chemical composition gradient are also considered. The result for the overshooting distance is l_{over}^{SMC} = 0.40_{−0.06}^{+0.12} H_{p} (Z_{0} = 0.004) and l_{over}^{LMC} = 0.10_{−0.10}^{+0.17} H_{p} (Z_{0} = 0.008) suggesting a possible dependence of the extent of the mixed central regions with metallicity within the considered mass range. Unfortunately it is not yet possible to fully disentangle effects of mass and chemical composition.

Accepted by: Astronomy & Astrophysics

For preprints, contact daniel.cordier@ensc-rennes.fr
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0206128

A statistical study of binary and multiple clusters in the LMC

A. Dieball^{1}, H. Müller^{2}, and E. K. Grebel^{3}

1 Sternwarte der Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2 Institut für Photogrammetrie, Universität Bonn, Nußallee 15, 53115 Bonn, Germany 3 MPI für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany

Based on the Bica et al. (1999, AJ, 117, 238) catalogue, we studied the star cluster system of the LMC and provide a new catalogue of all binary and multiple cluster candidates found. As a selection criterion we used a maximum separation of 1.′4 corresponding to 20 pc (assuming a distance modulus of 18.5 mag). We performed Monte Carlo simulations and produced artificial cluster distributions that we compared with the real one in order to check how many of the found cluster pairs and groups can be expected statistically due to chance superposition on the plane of the sky. We found that, depending on the cluster density, between 56% (bar region) and 12% (outer LMC) of the detected pairs can be explained statistically. We studied in detail the properties of the multiple cluster candidates. The binary cluster candidates seem to show a tendency to form with components of similar size. When possible, we studied the age structure of the cluster groups and found that the multiple clusters are predominantly young with only a few cluster groups older than 300 Myr. The spatial distribution of the cluster pairs and groups coincides with the distribution of clusters in general; however, old groups or groups with large internal age differences are mainly located in the densely populated bar region. Thus, they can easily be explained as chance superpositions. Our findings show that a formation scenario through tidal capture, as proposed by Vallenari et al. (1998, A&A, 331, 506), is not only unlikely due to the low probability of close encounters of star clusters, and thus the even lower probability of tidal capture, but the few groups with large internal age differences can easily be explained with projection effects. We favour a formation scenario as suggested by Fujimoto & Kumai (1997, AJ, 113,

249) in which the components of a binary cluster formed together and thus should be coeval or have small age differences compatible with cluster formation time scales.

Accepted by: A&A

For preprints, contact adieball@astro.uni-bonn.de

The Local Stellar Population of Nova regions in the Large Magellanic Cloud

Annapurni Subramaniam^{1} and G.C.Anupama^{1}

1 Indian Institute of Astrophysics, Bangalore, India

This study aims at identifying and understanding the parent population of novae in the Large Magellanic Cloud (LMC) by studying the local, projected, stellar population. The star formation history of the local environment around novae is studied based on photometric data of stars and star clusters in the nova neighbourhood, available in the OGLE II survey and star cluster catalogues. The age of the stellar population within a few arcmin around novae regions are estimated using isochrone fits to the V vs (V−I) colour-magnitude diagrams. The fraction of stars in various evolutionary states are compared using luminosity functions of the main-sequence stars and the red giant stars.

The age, density and luminosity function of the stellar population are estimated around 15 novae. The upper limit of the age of the intermediate stellar population is found to be 4 Gyr in all the regions, excepting the region around the slow nova LMC 1948. Star formation in these regions is found to have started between 4 – 2.0 Gyr, with a majority of the regions starting star formation at 3.2 Gyr. This star formation event lasted until 1.6 – 0.8 Gyr ago.

The star formation history of the underlying population of both the fast and moderately fast novae indicate their parent population to be similar and likely to be in the age range 3.2 – 1.0 Gyr. This is in good agreement with the theoretical age estimates for Galactic cataclysmic variables.

The region around the slow nova shows a stellar population in the age range 1 – 10 Gyr, with a good fraction older than 4 Gyr. This indicates that the progenitor might belong to an older population, consistent with the idea that the progenitors of slow novae belong to older population.

Accepted by: Astronomy & Astrophysics For preprints, contact purni@iiap.ernet.in Also available from astroph - 0203098

Mass segregation in rich clusters in the Large Magellanic Cloud III. Implications for the initial mass function

R. de Grijs,^{1} G.F. Gilmore,^{1} A.D. Mackey,^{1} M.I. Wilkinson,^{1} S.F. Beaulieu,^{2} R.A. Johnson,^{3} and B.X. Santiago^{4}

1 Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 2 Department of Physics and Astronomy, University of Victoria, 3800 Finnerty Road, Victoria, BC, V8P 1A1, Canada 3 European Southern Observatory, Casilla 19001, Santiago 19, Chile 4 Universidade Federal do Rio Grande do Sul, Instituto de Física, 91501-970 Porto Alegre, RS Brazil

The distribution of core radii of rich clusters in both the Large and Small Magellanic Clouds (LMC, SMC) systematically increases in both upper limit and spread with increasing cluster age. Clusterto-cluster variations in the stellar initial mass function (IMF) have been suggested as an explanation. We discuss the implications of the observed degree of mass segregation for the shape of the initial mass function.

Our results are based on Hubble Space Telescope/WFPC2 observations of six rich star clusters in the LMC, selected to include three pairs of clusters of similar age, mass, metallicity, and distance from the LMC centre, exhibiting a large spread in core radii between the clusters in each pair.

All clusters show clear evidence of mass segregation: (i) their luminosity function slopes steepen with increasing cluster radius, and (ii) the brighter stars are characterized by smaller core radii. For all sample clusters, both the slope of the luminosity function in the cluster centres and the degree of mass segregation are similar to each other, within observational errors of a few tenths of power-law slope fits to the data. This implies that their initial mass functions must have been very similar, down to ∼ 0.8 − 1.0M_{⊙}.

We rule out variations in the IMF as the main driver of the increasing spread of cluster core radii with cluster age. We propose instead that external tidal effects generate the range of core properties seen.

Submitted to: Monthly Notices of the Royal Astronomical Society For preprints, contact grijs@ast.cam.ac.uk

Fundamental Properties and Distances of LMC Eclipsing Binaries: III. EROS 1044

Ignasi Ribas^{1,2}, Edward L. Fitzpatrick^{1}, Frank P. Maloney^{1}, Edward F. Guinan^{1}, Andrzej Udalski^{3}

1 Department of Astronomy & Astrophysics, Villanova University, Villanova, PA 19085, USA 2 Departament d’Astronomia i Meteorologia, Universitat de Barcelona, Av. Diagonal, 647, E-08028 Barcelona, Spain 3 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present results from a detailed analysis of a third eclipsing binary (EB) system in the Large Magellanic Cloud, EROS 1044 (∼B2 IV–V + ∼B2 III–IV). Our study combines the “classical” EB study of light and radial velocity curves with detailed modeling of the observed spectral energy distribution, and yields an essentially complete picture of the stellar properties of the system and a determination of its distance. The observational data exploited include optical photometry, spacebased UV spectroscopy, and UV/optical spectrophotometry. The advantages of our technique include numerous consistency checks and, in the case of the distance determinations, the absence of zero point uncertainties and adjustable parameters. We find the EROS 1044 system to consist of a pair of normal, mildly-evolved ∼ 21000 K stars, whose derived properties are consistent with stellar evolution calculations. The distance to the system is 47.5 ± 1.8 kpc. We discuss the implications of our results for three EB systems (HV 2274, HV 982, and EROS 1044) on the general distance to the Large Magellanic Cloud.

Accepted by: The Astrophysical Journal For preprints, contact iribas@am.ub.es

Also available from the URL http://arXiv.org/abs/astro-ph/0204061

New Understanding of Large Magellanic Cloud Structure, Dynamics and Orbit from Carbon Star Kinematics

Roeland P. van der Marel^{1}, David R. Alves^{2}, Eduardo Hardy^{3} and Nicholas B. Suntzeff^{4}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 Columbia Astrophysics Laboratory, New York, NY 10027, USA 3 National Radio Astronomy Observatory, Casilla 36-D, Santiago, Chile 4 Cerro-Tololo Inter-American Observatory, Casilla 603, La Serena, Chile

We formulate a new, revised and coherent understanding of the structure and dynamics of the Large Magellanic Cloud (LMC), and its orbit around and interaction with the Milky Way. Much of our understanding of these issues hinges on studies of the LMC line-of-sight kinematics. The observed velocity field includes contributions from the LMC rotation curve V (R′), the LMC transverse velocity vector v⃗_{t}, and the rate of inclination change di/dt. All previous studies have assumed di/dt = 0. We show that this is incorrect, and that combined with uncertainties in v⃗_{t} this has led to incorrect estimates of many important structural parameters of the LMC. We derive general expressions for the velocity field which we fit to kinematical data for 1041 carbon stars. We calculate v⃗_{t} by compiling and improving LMC proper motion measurements from the literature, and we show that for known v⃗_{t} all other model parameters are uniquely determined by the data. The position angle of the line of nodes is Θ = 129.9^{∘} ± 6.0^{∘}, consistent with the value determined geometrically by van der Marel & Cioni (2001). The rate of inclination change is di/dt = −0.37±0.22mas yr^{−1} = −103±61 degrees/Gyr. This is similar in magnitude to predictions from N-body simulations by Weinberg (2000), which predict LMC disk precession and nutation due to Milky Way tidal torques. The LMC rotation curve V (R′) has amplitude 49.8 ± 15.9 km s^{−1}. This is 40% lower than what has previously (and incorrectly) been inferred from studies of HI, carbon stars, and other tracers. The line-of-sight velocity dispersion has an average value σ = 20.2 ± 0.5 km s^{−1}, with little variation as function of radius. The dynamical center of the carbon stars is consistent with the center of the bar and the center of the outer isophotes, but it is offset by 1.2^{∘} ± 0.6^{∘} from the kinematical center of the HI. The enclosed mass inside the last data point is M_{LMC}(8.9 kpc) = (8.7 ± 4.3) × 10^{9} M_{⊙}, more than half of which is due to a dark halo. The LMC has a considerable vertical thickness; its V/σ = 2.9 ± 0.9 is less than the value for the Milky Way’s thick disk (V/σ ≈ 3.9). Simple arguments for models stratified on spheroids indicate that the (out-of-plane) axial ratio could be ∼ 0.3 or larger. Isothermal disk models for the observed velocity dispersion profile confirm the finding of Alves & Nelson (2000) that the scale height must increase with radius. A substantial thickness for the LMC disk is consistent with the simulations of Weinberg, which predict LMC disk thickening due to Milky Way tidal forces. These affect LMC structure even inside the LMC tidal radius, which we calculate to be r_{t} = 15.0 ± 4.5 kpc (i.e., 17.1^{∘} ± 5.1^{∘}). The new insights into LMC structure need not significantly alter existing predictions for the LMC selflensing optical depth, which to lowest order depends only on σ. The compiled proper motion data imply an LMC transverse velocity v_{t} = 406 km s^{−1} in the direction of position angle 78.7^{∘} (with errors of ∼ 40 km s^{−1} in each coordinate). This can be combined with the observed systemic velocity, v_{sys} = 262.2 ± 3.4 km s^{−1}, to calculate the LMC velocity in the Galactocentric rest frame. This yields v_{LMC} = 293 ± 39 km s^{−1}, with radial and tangential components v_{LMC,rad} = 84 ± 7 km s^{−1} and v_{LMC,tan} = 281 ± 41 km s^{−1}, respectively. This is consistent with the range of velocities that has been predicted by models for the Magellanic Stream. The implied orbit of the LMC has an apocenter to pericenter distance ratio ∼ 2.5 : 1, a perigalactic distance ∼ 45 kpc, and a present orbital period around the Milky Way ∼ 1.5Gyr. The constraint that the LMC is bound to the Milky Way provides a robust limit on the minimum mass and extent of the Milky Way dark halo: M_{MW} ≥ 4.3×10^{11} M_{⊙} and r_{h} ≥ 39 kpc (68.3% confidence). Finally, we present predictions for the LMC proper motion velocity field, and we discuss how measurements of this may lead to kinematical distance estimates of the LMC.

Submitted to: The Astronomical Journal For preprints, contact marel@stsci.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0205161

ROSAT X-ray sources in the field of the LMC II.Statistics of background AGN and X-ray binaries

P. Kahabka^{1}

1 Sternwarte, Universität Bonn, Auf dem Hügel 71, D–53121 Bonn, Germany

About 200 X-ray sources from a sample of spectrally hard ROSAT PSPC sources, given in the catalog of Haberl & Pietsch (1999), and observed in a ∼60 square degree field of the LMC during several archival pointed observations with a wide range of exposure times have been reanalyzed. For these sources accurate count rates and hardness ratios have been recalculated. In comparison to Haberl & Pietsch (1999) we used merged data from all available observations and we derived average source parameters by investigating each source individually. From a simulation powerlaw spectral tracks have been derived in the HR1 – HR2 plane and ∼170 sources have been classified as background X-ray sources or as LMC X-ray binaries. 80% of the spectrally hard X-ray sources with more than 50 observed counts have been found to be consistent with background X-ray sources and 20% with LMC X-ray binaries (53 sources with AGN and 15 with X-ray binaries). The discovery of a new supersoft source RX J0529.4-6713 at the southern H i boundary of the supergiant shell LMC 4 is reported. We find two new candidate X-ray binary systems which are associated with the optical bar of the LMC and additional candidate X-ray binaries which are associated with supergiant shells.

Accepted by: Astronomy & Astrophysics

For preprints, contact pkahabka@astro.uni-bonn.de
Also available from the URL http://ArXiv.org/abs/astro-ph/0204076

ROSAT X-ray sources in the field of the LMC III. The log N – log S of background AGN and the LMC gas

P. Kahabka^{1}, K.S. de Boer^{1} and C. Brüns^{2}

1 Sternwarte, Universität Bonn, Auf dem Hügel 71, D–53121 Bonn, Germany 2 Radioastronomisches Institut, Universität Bonn, Auf dem Hügel 71, D–53121 Bonn, Germany

We use a sample of 50 background X-ray sources (AGN) and candidate AGN in the field of the LMC observed with more than 50 counts in archival ROSAT PSPC observations to derive the observed log N − log S relation. We correct for the inhomogenous ROSAT PSPC exposure and the varying absorption due to the galactic and the LMC gas (for which we used an H i map derived from observations with the Parkes radio telescope). We compare the observed log N − log S relation with a theoretical log N − log S relation of the soft extragalactic X-ray background (SXRB) which comprises an AGN and a cluster of galaxy contribution. We find that the observed log N − log S has a deficiency with respect to the theoretical log N − log S. There are several factors which can account for such a deficiency: (1) incompleteness of the selected AGN and cluster of galaxies sample, (2) deviation of the theoretical log N − log S in the LMC field from the log N − log S derived from a large sample of AGN in several fields in the sky, (3) the existence of gas additional to the H i represented in the Parkes H i map of the LMC field, restricted to the high column ≥ 10^{21} cm^{−2} regime. We investigate the likely contribution of these effects and find that (1) a fraction (of at most ∼30%) of the AGN and clusters of galaxies in the LMC field may not have been found in our analysis and may contribute to the observed deficiency. The existence of extended regions with hot diffuse gas and source crowding makes the detection of all AGN and clusters of galaxies very difficult. (2) We cannot exclude a deviation of the log N − log S in the field of the LMC from a mean theoretical log N − log S, especially the cluster of galaxy contribution which is of importance in the flux range we are comprising may show variations across the sky. (3) If LMC gas in addition to the H i represented in the Parkes H i map would be responsible for the deficiency and if this additional gas is restricted to the high column ≥ 10^{21} cm^{−2} regime, and assuming that the metallicity of the ISM of the LMC is −0.3 dex lower than the metallicity of the galactic ISM, then a factor of 1.9±_{1.6}^{3.3} at 90% confidence of additional gas would be required which, if purely molecular, would be equal to a molecular mass fraction of 63 ±_{42}^{20} %. Such a value would be larger than but within the uncertainties consistent with a molecular mass fraction of ∼30% derived from CO observations for the high column regime of the LMC gas. From this analysis, it follows that some gas additional to the measured H i for the high column regime of the LMC gas is likely to be required to explain the observed log N − log S. But the amount of such additional gas is dependent on the completeness of our selected AGN and clusters of galaxies sample and on the assumptions made about the description of the log N − log S of the SXRB in the field of the LMC.

Accepted by: Astronomy & Astrophysics

For preprints, contact pkahabka@astro.uni-bonn.de
Also available from the URL http://ArXiv.org/abs/astro-ph/0204096

Contribution of the Large Magellanic Cloud to the Galactic Warp

T. Tsuchiya^{1}

1 Astronomisches Rechen-Institute, Mönchhofstraße 12–14, D69120 Heidelberg, Germany

Multi-scale interaction between the LMC, Galactic halo, and the disk is examined with N-body simulations, and precise amplitudes of the Galactic warp excitation are obtained. The Galactic models are constructed most realistically to satisfy available observational constraints on the local circular velocity, the mass, surface density and thickness of the disk, the mass and size of the bulge, the local density of the halo matter at the solar radius, and the mass and orbit of the LMC. The mass of the halo within R = 50 kpc is set to about 5 × 10^{11}M_{⊙}. Since observational estimate of the mass distributed in outer region has large ambiguity, two extreme cases are examined; M(< 170kpc) = 2.1 and 0.9 × 10^{12}M_{⊙}. LMC is orbiting in a ellipse with apocentric radii of 100 kpc, thus the main difference between our two models is the mass density in the satellite orbiting region, so that our study can clarify the role of the halo on excitation of the warp.

By using hybrid algorithm (SCF-TREE) I have succeeded to follow the evolution with millions of particles. The orbiting satellite excites density enhancement as a wake, and the wake exerts a tidal force on the disk. Because of the additional torque from the wakes in the halo, the amplitudes of the induced warps are much larger than the classical estimate by Hunter & Toomre (1969), who considered only the direct torque from the LMC. The obtained amplitudes of m = 0, 1, 2 warps in the larger halo model show very good agreement with the observed amplitude in the Milky Way. This result revives the LMC as a possible candidate of the origin of the Galactic warp. Our smaller halo model, however, yield only weak warps in all the harmonic modes. Therefore the halo still has significant influence on excitation of warp even in the interaction scenario for excitation of warps.

Accepted by: New Astronomy

For preprints, contact tsuchiya@ari.uni-heidelberg.de

Discovery of four X-ray quasars behind the Large Magellanic Cloud

A. Dobrzycki^{1}, P. J. Groot^{1,2}, L. M. Macri^{1} and K. Z. Stanek^{1}

1 Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge MA 02138, USA 2 Present address: Department of Astrophysics, University of Nijmegen, PO Box 9010, 6500 GL Nijmegen, The Netherlands

We present the discovery of four X-ray quasars (z_{em} = 0.26, 0.53, 0.61, 1.63) located behind the Large Magellanic Cloud; three of them are located behind the bar of the LMC. The quasars were identified via spectroscopy of optical counterparts to X-ray sources found serendipitously by the Chandra X-ray Observatory satellite. All four quasars have archival V I photometry from the OGLE-II project; one of them was found by OGLE to be variable. We present the properties of the quasars and discuss their possible applications.

Submitted to: Ap.J.Letters

For preprints, contact adobrzycki@cfa.harvard.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0202524

Abstracts of Non-Refereed Papers

First deep XMM-Newton observations of the LMC: Identifying LMC intrinsic source populations

F. Haberl

Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße 1, 85748 Garching, Germany

The first X-ray survey of the Large Magellanic Cloud (LMC) was performed with the Einstein satellite, revealing diffuse X-ray emission from hot gas and discrete X-ray sources. The ROSAT observations between 1990 and 1998 supplied the most sensitive survey with imaging instruments in the soft X-ray band (0.1 - 2.4 keV). The PSPC and HRI observations covered 59 square degrees of the LMC and yielded a catalogue of about 1000 sources.

Large efforts were undertaken to identify and classify the X-ray sources according to the nature of their X-ray emission. X-ray properties were used together with information from other electromagnetic wavelength bands to screen out foreground stars and background objects from the LMC intrinsic X-ray source population which comprises X-ray binaries, supernova remnants and supersoft sources. However, the vast majority of sources still remains of unknown nature.

First deep XMM-Newton observations of selected regions in the LMC demonstrate a large step forward in the identification of LMC X-ray sources. The large collecting area together with imaging detectors covering energies 0.1 - 15 keV with far improved spectral resolution allows to determine the nature of an object directly from the X-ray properties down to low flux levels of a few 10^{−14} erg cm^{−2} s^{−1}. First results of a deep observation north of LMC X-4 are presented, which reveal the presence of new supernova remnants and X-ray binaries.

To appear in: Proceedings of the Symposium on ‘New Visions of the X-ray Universe in the XMM-Newton and Chandra Era’, 26-30 November 2001, ESTEC, The Netherlands For preprints, contact fwh@mpe.mpg.de

Also available from the URL ftp://ftp.xray.mpe.mpg.de/people/fwh/docs/fhaberl-E3.ps.gz

The population of X-ray binaries and background AGN in the field of the Large Magellanic Cloud

P. Kahabka^{1}

1 Sternwarte, Universität Bonn, Auf dem Hügel 71, D–53121 Bonn, Germany

A sample of ∼30 (candidate) X-ray binaries and ∼140 (candidate) background AGN in the LMC field was derived from the ROSAT PSPC survey and has been discussed in Kahabka (2002, A&A 388, 100). A log N − log S analysis of background sources in the LMC field with more than 50 observed counts has been performed by Kahabka et al. (2002, A&A 388, 113). Here a summary of the X-ray binary and background AGN sample is given taking a few new identifications into account. The catalog of 50 X-ray sources with unabsorbed fluxes (0.5 – 2.0 keV) in excess of 1.4 10^{−14} erg cm^{−2} s^{−1} and used for the log N − log S analysis in the LMC field is given. Consideration of obscuration by dark absorbing clouds of 15% of the LMC field used for the log N − log S analysis gives consistency between the observed log N − log S and the theoretical log N − log S of background sources (Gilli et al. 2001, A&A 366, 407). The estimated mass associated with these clouds is ∼5 10^{7} M_{⊙}, consistent with the total mass of molecular clouds in the LMC inferred from CO surveys.

To appear in: “XEUS – studying the evolution of the hot universe”, Conference Proceedings held in Garching, Germany, March 11–15, 2002, eds. G. Hasinger, T. Boller and A. Parmar

For preprints, contact pkahabka@astro.uni-bonn.de

Thesis Abstract

Untersuchungen zur Entwicklung von Satellitengalaxien

Björn Seidel

Max–Planck–Institut für Astronomie; Königstuhl 17; D–69117 Heidelberg; Germany

Equilibrium models of one- and two-component satellite galaxies are created. The tidal influence of the Milky Way, which is modeled as a rigid external potential, is examined.

A first set of simulations adopting originally spherically-symmetric, one-component satellites show that after elliptical deformation, a bar and tails of unequal length arise which change their morphology cyclically. By considering comparative simulations, the following phenomena are discovered: (1) High– density regions in the tails, (2) low–density zones around the core or bar, (3) an often concealed bar. The overall morphology as function of time is analysed. The particles are lost from the core via the bar and move along certain morphologically characteristic structures into the tails.

After deducing general quantities of the multi-component King model, three stable standard models of two-component satellite galaxies with different distributions of dark and visible matter are found. These models have a mass ratio of 1:10 for baryonic to dark matter. Without restriction to the universality of the results, the Large Magellanic Cloud was taken as basic prototype for these models. To choose the proper orbits, the tidal radius of the three-component model of the potential of the Galaxy is calculated both analytically and numerically. Then the simulations are analysed with the different behaviour of the components being the main focus of interest.

Some of the questions addressed and answered here are: Is it possible to detach a large fraction of dark matter but only a small amount of visible matter? Do dark and visible matter form different morphological structures? What determines the morphology of the components? How do the dissolution and the decay of the core of a satellite occur in detail? How do the properties of the core (densities, velocity dispersions,...) depend on time? How is the radial dependence of the velocity dispersion? How exactly does the satellite galaxy decay?

PhD Thesis completed at the Ruprecht–Karls–Universität Heidelberg on January 30, 2002 under the direction of Dr. Andreas Burkert (MPIA).

Available from the URLs http://www.diplomaxx.de/katalog/dissertation/...
.../dis-astronomie-theoretische as/0238/body 0238.html
and http://www.tk-diplom.de/Katalog/A1b/A1bL/a1bl 14.html .

For more information, contact seidel@mpia-hd.mpg.de